# Hybrid Feedback for Three-dimensional Convex Obstacle Avoidance (Extended version)

### v2

@@ IX-E Proof of Lemma 4
First, we prove that when the control input corresponds to the obstacle-avoidance mode, $\mathbf{x}(t,j)\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$ for all $(t,j)\in(I_{j_{1}+1},j_{1}+1)$. To that end, we make use of Nagumo’s theorem [8, Theorem 4.7] and show that when the control input corresponds to the obstacle-avoidance mode, one has
@@
for all $\mathbf{x}\in\partial\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$. This, combined with the fact that the control vector trajectory $\mathbf{u}(\xi)$ is continuous, when it corresponds to the obstacle-avoidance mode, as stated in Lemma 2, ensures that $\mathbf{x}(t,j)\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$ for all $(t,j)\in(I_{j_{1}+1}\times j_{1}+1)$.
Note that $\partial\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))=\partial\mathcal{D}_{r_{a}}(\mathcal{O}_{i})\cup\partial\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{i})$. For all $\mathbf{x}\in\partial\mathcal{D}_{r_{a}}(\mathcal{O}_{i})$, one has $\mathcal{P}_{\geq}(\mathbf{0},\mathbf{x}_{\pi})\subset\mathbf{T}_{\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))}(\mathbf{x})$, where $\mathbf{x}_{\pi}=\mathbf{x}-\Pi(\mathbf{x},\mathcal{O}_{\mathcal{W}})$ with $\Pi(\mathbf{x},\mathcal{O}_{\mathcal{W}})=\Pi(\mathbf{x},\mathcal{O}_{i})$. Also, since the control input corresponds to the obstacle-avoidance mode, for $\mathbf{x}\in\partial\mathcal{D}_{r_{a}}(\mathcal{O}_{i})$, the control vector (12a) is given by $\mathbf{u}(\xi)=\kappa_{r}\mathbf{P}(\mathbf{a})\mathbf{x}_{\pi}$. Now, for any $\mathbf{p}\in\mathbb{R}^{3}$ and a unit vector $\mathbf{q}\in\mathbb{S}^{2}$, it can be shown that $\mathbf{p}^{\top}\mathbf{P}(\mathbf{q})\mathbf{p}\geq 0$. Therefore, for all $\mathbf{x}\in\partial\mathcal{D}_{r_{a}}(\mathcal{O}_{i})$, one has $\mathbf{x}_{\pi}\mathbf{P}(\mathbf{a})\mathbf{x}_{\pi}\geq 0$. This implies that $\mathbf{u}(\xi)\in\mathcal{P}_{\geq}(\mathbf{0},\mathbf{x}_{\pi})\subset\mathbf{T}_{\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))}(\mathbf{x})$ for $\mathbf{x}\in\partial\mathcal{D}_{r_{a}}(\mathcal{O}_{i})$, and condition (45) holds true.
Next, for $\mathbf{x}\in\partial\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{i})$, one has $\mathcal{P}_{\leq}(\mathbf{0},\mathbf{x}_{\pi})\subset\mathbf{T}_{\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))}(\mathbf{x})$. Also, since the control input corresponds to the obstacle-avoidance mode, for $\mathbf{x}\in\partial\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{i})$, the control vector (12a) is given by $\mathbf{u}(\xi)=-\kappa_{r}\mathbf{P}(\mathbf{a})\mathbf{x}_{\pi}$. As mentioned earlier, for all $\mathbf{x}\in\partial\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{i})$, one has $\mathbf{x}_{\pi}\mathbf{P}(\mathbf{a})\mathbf{x}_{\pi}\geq 0$. Therefore, $\mathbf{u}(\xi)\in\mathcal{P}_{\leq}(\mathbf{0},\mathbf{x}_{\pi})\subset\mathbf{T}_{\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))}(\mathbf{x})$ for $\mathbf{x}\in\partial\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{i})$, and condition (45) holds true. As a result, since $\mathbf{x}(t_{1},j_{1}+1)\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$, for all $(t,j)\in(I_{j_{1}+1},j_{1}+1)$, one can conclude that
$\mathbf{x}(t,j)\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i})).$ (46)
@@
Let us define the set $\mathcal{O}_{i}^{\mathcal{S}}=\mathcal{D}_{r_{a}}(\mathcal{O}_{i})\cap\mathcal{P}(\mathbf{h},\mathbf{a})$, as shown in Fig. 5. Since obstacle $\mathcal{O}_{i}$ is convex, the set $\mathcal{O}_{i}^{\mathcal{S}}$ is also convex. As a result, the target location has a unique closest point on the set $\mathcal{O}_{i}^{\mathcal{S}}$, represented by $\Pi(\mathbf{0},\mathcal{O}_{i}^{\mathcal{S}}).$ Let us define the set $\mathcal{LS}:=\mathcal{L}_{s}(\mathbf{0},\Pi(\mathbf{0},\mathcal{O}_{i}^{\mathcal{S}}))\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$. According to Remark 3, one has $\mathbf{0}\notin\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{i})$. Therefore, it is clear that $\mathcal{LS}\cap\partial\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i})\neq\emptyset$ for all $\beta\in[0,\gamma]$. Since $\mathbf{0}\in\mathcal{P}(\mathbf{h},\mathbf{a}),$ the line segment $\mathcal{LS}$ belongs to the plane $\mathcal{P}(\mathbf{h},\mathbf{a}).$ Since $\mathcal{LS}\cap\mathcal{D}_{r_{a}}^{\circ}(\mathcal{O}_{i})=\emptyset$, the line segment $\mathcal{LS}$ also belongs to the exit region $\mathcal{R}_{e}$ (18). Since the hit point $\mathbf{h}$ belongs to $\mathcal{R}_{l}^{i}$, the target location $\mathbf{0}$ is closer to the location $\Pi(\mathbf{0},\mathcal{O}_{i}^{\mathcal{S}})$ than to the hit point $\mathbf{h}$. Hence, for a sufficiently small value of $\bar{\epsilon}$, used in (23), one can ensure that the set $\mathcal{LS}$ belongs to the set $\mathcal{J}_{1}^{\mathcal{W}}$ (22).
@@
First we show that when the control input corresponds to the obstacle-avoidance mode and the state $\mathbf{x}$ is either in the set $\mathcal{S}_{1}$ or in the set $\mathcal{S}_{3}$, the control will eventually steer the state $\mathbf{x}$ to the set $\mathcal{S}_{2}.$
@@
Fact 1: Let us consider a hyperplane $\mathcal{P}(\mathbf{p},\mathbf{q})$, where $\mathbf{p}\in\mathcal{J}_{0}^{\mathcal{W}}\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$, for some $i\in\mathbb{I}$ and $\mathbf{q}=\mathbf{A}(\mathbf{p})$, where the mapping $\mathbf{A}$ is defined in (29). Then, for all $\mathbf{x}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{p},\mathbf{q})$, one has $\mathbf{x}_{\pi}\mathbf{P}(\mathbf{q})\mathbf{x}_{\pi}>0$.
@@
This proof is by contradiction. First, note that for any vector $\mathbf{s}\in\mathbb{R}^{n}$, one has $\mathbf{s}^{\top}\mathbf{P}(\mathbf{q})\mathbf{s}\geq 0$. Let us assume that there exists $\mathbf{x}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{q})$ such that $\mathbf{x}_{\pi}\mathbf{P}(\mathbf{q})\mathbf{x}_{\pi}=0$. Since $\mathbf{x}_{\pi}\neq\mathbf{0}$, one has $\mathbf{P}(\mathbf{q})\mathbf{x}_{\pi}=\mathbf{0}$. Therefore, the vector $\mathbf{x}_{\pi}$ is normal to the hyperplane $\mathcal{P}(\mathbf{p},\mathbf{q})$. This implies that that plane $\mathcal{P}(\mathbf{p},\mathbf{q})$ is a supporting hyperplane [19, Section 2.5.2] to the convex set $\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i})$ at $\mathbf{x}$, where $\beta=d(\mathbf{x},\mathcal{O}_{i})-r_{a}\in[0,\gamma]$. Therefore, the set $\mathcal{D}_{r_{a}}^{\circ}(\mathcal{O}_{i})\cap\mathcal{P}(\mathbf{p},\mathbf{q})$ is an empty set.
However, since $\mathbf{q}=\mathbf{A}(\mathbf{p})$, according to (29), one has $\mathbf{p}_{\pi}\in\mathcal{P}(\mathbf{p},\mathbf{q})$. Therefore, $\mathcal{L}(\mathbf{p},\Pi(\mathbf{p},\mathcal{O}_{i}))\subset\mathcal{P}(\mathbf{p},\mathbf{q})$. As a result, $\mathcal{L}(\mathbf{p},\Pi(\mathbf{p},\mathcal{O}_{i}))\cap\mathcal{D}_{r_{a}}^{\circ}(\mathcal{O}_{i})\neq\emptyset$. This implies that $\mathcal{D}_{r_{a}}^{\circ}(\mathcal{O}_{i})\cap\mathcal{P}(\mathbf{p},\mathbf{q})\neq\emptyset$, which is a contradiction. ∎
According to Fact 1, for all $\mathbf{x}\in\partial\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i})\cap\mathcal{S}_{1}$, where $\beta\in[0,\gamma_{a}]$, one has $\mathbf{x}_{\pi}\mathbf{P}(\mathbf{a})\mathbf{x}_{\pi}>0$. Therefore, as discussed earlier, the control input vector (49) steers $\mathbf{x}$ to the interior of the set $(\mathcal{S}_{1}\cup\mathcal{S}_{2})\setminus\mathcal{D}_{r_{a}+\beta}^{\circ}(\mathcal{O}_{i})$ and eventually $\mathbf{x}$ will enter in the set $\mathcal{S}_{2}$.
@@
According to Fact 1, for all $\mathbf{x}\in\partial\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i})\cap\mathcal{S}_{3}$, where $\beta\in[\gamma_{s},\gamma]$, one has $\mathbf{x}_{\pi}\mathbf{P}(\mathbf{a})\mathbf{x}_{\pi}>0$. This implies that the control input vector (50) steers $\mathbf{x}$ to the interior of the set $\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i})$. This, combined with the fact that $\mathbf{x}(t,j)\in\mathcal{P}(\mathbf{h},\mathbf{a})$, for all $(t,j)\in(I_{j_{1}+1}\times j_{1}+1)$, as per claim 1 in Lemma 4, ensures that the control input vector (50) steers $\mathbf{x}$ to the interior of the set $(\mathcal{S}_{3}\cup\mathcal{S}_{2})\cap\mathcal{D}_{r_{a}+\beta}^{\circ}(\mathcal{O}_{i})$ and eventually $\mathbf{x}$ will enter in the set $\mathcal{S}_{2}$.

### v3

@@ IX-E Proof of Lemma 4
First, we prove that when the control input corresponds to the obstacle-avoidance mode, one has $\mathbf{x}(t,j)\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$ for all $(t,j)\in(I_{j_{1}+1},j_{1}+1)$. To that end, we make use of Nagumo’s theorem [11, Theorem 4.7] and show that when the control input corresponds to the obstacle-avoidance mode, one has
@@
for all $\mathbf{x}\in\partial\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$. This, combined with the fact that the control input vector $\mathbf{u}(\xi)$ is continuous, when it corresponds to the obstacle-avoidance mode, as stated in Lemma 2, ensures that $\mathbf{x}(t,j)\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$ for all $(t,j)\in(I_{j_{1}+1}\times j_{1}+1)$.
Note that $\partial\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))=\partial\mathcal{D}_{r_{a}}(\mathcal{O}_{i})\cup\partial\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{i})$. For all $\mathbf{x}\in\partial\mathcal{D}_{r_{a}}(\mathcal{O}_{i})$, one has $\mathcal{P}_{\geq}(\mathbf{0},\mathbf{x}_{\pi})\subset\mathbf{T}_{\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))}(\mathbf{x})$, where $\mathbf{x}_{\pi}=\mathbf{x}-\Pi(\mathbf{x},\mathcal{O}_{\mathcal{W}})$ with $\Pi(\mathbf{x},\mathcal{O}_{\mathcal{W}})=\Pi(\mathbf{x},\mathcal{O}_{i})$. Also, since the control input corresponds to the obstacle-avoidance mode, for $\mathbf{x}\in\partial\mathcal{D}_{r_{a}}(\mathcal{O}_{i})$, the control vector (11a) is given by $\mathbf{u}(\xi)=\kappa_{r}\mathbf{P}(\mathbf{a})\mathbf{x}_{\pi}$. Since for any $\mathbf{q}\in\mathbb{S}^{2}$, the matrix $\mathbf{P}(\mathbf{q})$ is positive semidefinite, one has $\mathbf{p}^{\top}\mathbf{P}(\mathbf{q})\mathbf{p}\geq 0$ for all $\mathbf{p}\in\mathbb{R}^{3}$. Therefore, for all $\mathbf{x}\in\partial\mathcal{D}_{r_{a}}(\mathcal{O}_{i})$, it is true that $\mathbf{x}_{\pi}\mathbf{P}(\mathbf{a})\mathbf{x}_{\pi}\geq 0$. This implies that $\mathbf{u}(\xi)\in\mathcal{P}_{\geq}(\mathbf{0},\mathbf{x}_{\pi})\subset\mathbf{T}_{\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))}(\mathbf{x})$ for $\mathbf{x}\in\partial\mathcal{D}_{r_{a}}(\mathcal{O}_{i})$, and condition (44) holds true.
Next, for $\mathbf{x}\in\partial\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{i})$, one has $\mathcal{P}_{\leq}(\mathbf{0},\mathbf{x}_{\pi})\subset\mathbf{T}_{\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))}(\mathbf{x})$. Also, since the control input corresponds to the obstacle-avoidance mode, for $\mathbf{x}\in\partial\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{i})$, the control vector (11a) is given by $\mathbf{u}(\xi)=-\kappa_{r}\mathbf{P}(\mathbf{a})\mathbf{x}_{\pi}$. As mentioned earlier, for all $\mathbf{x}\in\partial\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{i})$, one has $\mathbf{x}_{\pi}\mathbf{P}(\mathbf{a})\mathbf{x}_{\pi}\geq 0$. Therefore, $\mathbf{u}(\xi)\in\mathcal{P}_{\leq}(\mathbf{0},\mathbf{x}_{\pi})\subset\mathbf{T}_{\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))}(\mathbf{x})$ for $\mathbf{x}\in\partial\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{i})$, and condition (44) holds true. As a result, since $\mathbf{x}(t_{1},j_{1}+1)\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$, one can conclude that
$\mathbf{x}(t,j)\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i})),$ (45)
for all $(t,j)\in(I_{j_{1}+1},j_{1}+1)$.
@@
Let us define the set $\mathcal{O}_{i}^{\mathcal{S}}=\mathcal{D}_{r_{a}}(\mathcal{O}_{i})\cap\mathcal{P}(\mathbf{h},\mathbf{a})$, as shown in Fig. 6. Since obstacle $\mathcal{O}_{i}$ is convex, the set $\mathcal{O}_{i}^{\mathcal{S}}$ is also convex. As a result, the target location has a unique closest point on the set $\mathcal{O}_{i}^{\mathcal{S}}$, represented by $\Pi(\mathbf{0},\mathcal{O}_{i}^{\mathcal{S}}).$ We define a set $\mathcal{LS}$ as follows:
$\mathcal{LS}:=\mathcal{L}(\mathbf{0},\mathbf{0}_{\pi}^{\mathcal{S}})\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}_{\geq}(\mathbf{0}_{\pi}^{\mathcal{S}},-\mathbf{0}_{\pi}^{\mathcal{S}}),$ (46)
where $\mathbf{0}_{\pi}^{\mathcal{S}}=\Pi(\mathbf{0},\mathcal{O}_{i}^{\mathcal{S}})$. Since $\mathbf{0}\in\mathcal{P}(\mathbf{h},\mathbf{a})$ and $\mathcal{O}_{i}^{\mathcal{S}}\subset\mathcal{P}(\mathbf{h},\mathbf{a})$, the line segment $\mathcal{LS}$ belongs to the plane $\mathcal{P}(\mathbf{h},\mathbf{a}).$ Since $\mathcal{LS}\cap\mathcal{D}_{r_{a}}^{\circ}(\mathcal{O}_{i})=\emptyset$, the line segment $\mathcal{LS}$ also belongs to the exit region $\mathcal{R}_{e}$ (17). Since the hit point $\mathbf{h}$ belongs to $\mathcal{R}_{l}^{i}$, the target location $\mathbf{0}$ is closer to $\Pi(\mathbf{0},\mathcal{O}_{i}^{\mathcal{S}})$ than to $\mathbf{h}$. Hence, if $\mathbf{0}\notin\mathcal{D}_{r_{a}+\gamma}(\mathcal{O}_{i})$, then for a sufficiently small value of $\bar{\epsilon}$, used in (22), one can ensure that the set $\mathcal{LS}$ belongs to the set $\mathcal{J}_{1}^{\mathcal{W}}$ (21). On the other hand, if $\mathbf{0}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$, it is straightforward to verify that $\mathcal{LS}\subset\mathcal{S}_{0}$, which, according to (21), implies that $\mathcal{LS}\subset\mathcal{J}_{1}^{\mathcal{W}}$.
@@
First, we show that when the control input corresponds to the obstacle-avoidance mode and the state $\mathbf{x}$ is either in the set $\mathcal{S}_{1}$ or in the set $\mathcal{S}_{3}$, the control will eventually steer the state $\mathbf{x}$ to the set $\mathcal{S}_{2}.$
@@
Fact 1: Let us consider the hyperplane $\mathcal{P}(\mathbf{p},\mathbf{q})$, where $\mathbf{p}\in\mathcal{J}_{0}^{\mathcal{W}}\cap\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))$, for some $i\in\mathbb{I}$ and $\mathbf{q}\in\mathbf{A}(\mathbf{p})$, where the mapping $\mathbf{A}$ is defined in (28). Then, for all $\mathbf{x}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{p},\mathbf{q})$, one has $\mathbf{x}_{\pi}^{\top}\mathbf{P}(\mathbf{q})\mathbf{x}_{\pi}>0$.
@@
This proof is by contradiction. First, note that for any $\mathbf{q}\in\mathbb{S}^{2}$, the matrix $\mathbf{P}(\mathbf{q})$ is positive semidefinite. Therefore, for any vector $\mathbf{s}\in\mathbb{R}^{n}$, one has $\mathbf{s}^{\top}\mathbf{P}(\mathbf{q})\mathbf{s}\geq 0$. Let us assume that there exists $\mathbf{x}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{p},\mathbf{q})$ such that $\mathbf{x}_{\pi}^{\top}\mathbf{P}(\mathbf{q})\mathbf{x}_{\pi}=0$. Since $\mathbf{x}_{\pi}\neq\mathbf{0}$, one has $\mathbf{P}(\mathbf{q})\mathbf{x}_{\pi}=\mathbf{0}$. Therefore, the vector $\mathbf{x}_{\pi}$ is normal to the hyperplane $\mathcal{P}(\mathbf{p},\mathbf{q})$. This implies that the plane $\mathcal{P}(\mathbf{p},\mathbf{q})$ is a supporting hyperplane [26, Section 2.5.2] to the convex set $\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i})$ at $\mathbf{x}$, where $\beta=d(\mathbf{x},\mathcal{O}_{i})-r_{a}\in[0,\gamma]$. Therefore, the set $\mathcal{D}_{r_{a}}^{\circ}(\mathcal{O}_{i})\cap\mathcal{P}(\mathbf{p},\mathbf{q})$ is an empty set.
However, since $\mathbf{q}\in\mathbf{A}(\mathbf{p})$, according to (28), one has $\mathbf{p}_{\pi}\in\mathcal{P}(\mathbf{p},\mathbf{q})$ and $\mathbf{0}\in\mathcal{P}(\mathbf{p},\mathbf{q})$. Therefore, $\mathcal{L}(\mathbf{p},\Pi(\mathbf{p},\mathcal{O}_{i}))\subset\mathcal{P}(\mathbf{p},\mathbf{q})$. As a result, $\mathcal{L}(\mathbf{p},\Pi(\mathbf{p},\mathcal{O}_{i}))\cap\mathcal{D}_{r_{a}}^{\circ}(\mathcal{O}_{i})\neq\emptyset$. This implies that $\mathcal{D}_{r_{a}}^{\circ}(\mathcal{O}_{i})\cap\mathcal{P}(\mathbf{p},\mathbf{q})\neq\emptyset$, which is a contradiction. ∎
According to Fact 1, for all $\mathbf{x}\in\partial\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i})\cap\mathcal{S}_{1}$, where $\beta\in[0,\gamma_{a}]$, one has $\mathbf{x}_{\pi}^{\top}\mathbf{P}(\mathbf{a})\mathbf{x}_{\pi}>0$. Therefore, as discussed earlier, the control input vector (49) steers $\mathbf{x}$ to the interior of the set $(\mathcal{S}_{1}\cup\mathcal{S}_{2})\setminus\mathcal{D}_{r_{a}+\beta}^{\circ}(\mathcal{O}_{i})$ and eventually $\mathbf{x}$ will enter in the set $\mathcal{S}_{2}$.
@@
According to Fact 1, for all $\mathbf{x}\in\partial\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i})\cap\mathcal{S}_{3}$, where $\beta\in[\gamma_{s},\gamma]$, one has $\mathbf{x}_{\pi}^{\top}\mathbf{P}(\mathbf{a})\mathbf{x}_{\pi}>0$. This implies that the control input vector (50) steers $\mathbf{x}$ to the interior of the set $\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i})$. This, combined with the fact that $\mathbf{x}(t,j)\in\mathcal{P}(\mathbf{h},\mathbf{a})$, for all $(t,j)\in(I_{j_{1}+1}\times j_{1}+1)$, as per claim 1 in Lemma 4, ensures that the control input vector (50) steers $\mathbf{x}$ to the interior of the set $(\mathcal{S}_{3}\cup\mathcal{S}_{2})\cap\mathcal{D}_{r_{a}+\beta}^{\circ}(\mathcal{O}_{i})$ and eventually $\mathbf{x}$ will enter in the set $\mathcal{S}_{2}$.